OpenStudy (anonymous):

If a pro basketball player has a vertical leap of about 20 inches, what is his hang time? Use the hang-time function V=48 T^2 His hang time is ? seconds. (Simplify your answer. Type an integer or a decimal. Round to the nearest tenth.)

OpenStudy (anonymous):

Given: V = 20 inches and V = 48 T^2 To find the time T, substitute V = 20 So it becomes 20 = 48 T^2 T^2 = 20/48 = 5/12 Taking square root on both sides , we get, T = Sqrt(5/12) T = 0.6457553 Rounded up to the nearest tenth, we got T = 0.6

OpenStudy (anonymous):

So the answer is 0.6 seconds?

OpenStudy (anonymous):

Got it!!! THX!! ^_^